The current asset section of the Excalibur Tire Company’s balance sheet consists of cash, marketable securities, accounts receivable, and inventories. The December 31, 2018, balance sheet revealed the following: Inventories $ 880,000 Total assets $ 3,200,000 Current ratio 2.50 Acid-test ratio 1.50 Debt to equity ratio 1.5 Required: Determine the following 2018 balance sheet items:

1. Current assets

2.Shareholder's equity

3. Long term assets

4.Long term liabilities
Answer #1

1) Acid test ratio = Current assets-Inventory/Current liabilities

1.50 = 2.50X-880000/X

1.50X = 2.50X-880000

1X = 880000

Current liabilities = 880000 and current assets = 880000*2.5 = 2200000

2) Total debt and equity = 3200000

Debt = 3200000*1.5/2.5 = 1920000

Equity 1920000/1.5 = 1280000

3) Long term assets = 3200000-2200000 = 1000000

4) Long term liabilities = 1920000-880000 = 1040000